Table 17.2 Lennard-Jones Parameters for Atoms and Molecules Particle e/kJ mol-1 o/Å 0.997 1.77 0.307 0.765 0.943 3.40 4.10 2.Using this table determine the van der Waal radius for Ar. Then use this radius to determine the fraction of volume occupied by 1 mol of argon at 25C and 1 atm.
Answer #1

The minimum van der Waals radius (r) can be calculated as follows.

r = \sigma/2

= (3.4/2) Ao

= 1.7 Ao

Volume = 4/3 \pir3 = 4/3 \pi(1.7*10-8 cm)3 = 20.5795*10-24 cm3 = 2.06*10-27 L (Note: 103 cm3 = 1 L)